what is the sum of the angles of a polygon with 22 sides?
Answer by venugopalramana(3286)

what is the sum of the angles of a polygon with 22 sides?
SUM OF ANGLES IN A POLYGON OF N SIDES IS GIVEN BY
S=(2N-4)*90..DEGREES
FOR N=22
S=(2*22-4)*90=18*90=1620 DEGREES